OpenStudy (anonymous):

Solve: 10t^2 - 9t = 1

OpenStudy (anonymous):

10t^2 - 9t - 1 = 0 10t^2 -10t + t -1=0 10t(t-1) +1(t-1)=0 (10t + 1)(t-1)=0 t = -1/10 and 1